# The Why Behind the Action: Unveiling Internal Drivers via Agentic Attribution

### v1

@@ 4.3 Quantitative Analysis
Experimental settings. We use the same model and cases as described in Section 4.2. For each case, five human annotators independently annotate the attribution ground truth, identifying the sentences that are most likely to contribute to the model’s final action. To improve annotation reliability, we take the intersection of the annotations provided by all five annotators as the final ground truth. All methods are implemented based on their respective official GitHub repositories (Liu et al., 2024; Cohen-Wang et al., 2024; Qi et al., 2024).
@@
Prob. Drop&Hold 0.944 1.000 1.000
Leave-one-out (LOO) 0.833 0.944 1.000
ContextCite 0.833 0.944 0.944
Saliency Score***Due to out-of-memory (OOM) issues in gradient-based salience computation on long cases, metrics are reported only for successfully completed cases. 0.700 1.000 1.000
@@ 5 Related Work
LLM-based Agent. In recent years, large language model (LLM)-based agents have emerged as a popular paradigm for human-AI collaboration and automated task execution, with two core modules that extend beyond single-turn context reasoning: memory and tool. Exsiting work on agent memory can be broadly catagorized into short-term (Zhong et al., 2024), long-term (Lewis et al., 2020) and structured (Packer et al., 2023) memory. By storing, retrieving, and updating salient information across interactions, they enable agents to maintain long-term context, user preferences, and goals, thereby supporting more coherent and personalized decision-making in multi-step tasks. Meanwhile, (Yao et al., 2022) (Schick et al., 2023) provide paradigms for enhancing agents’ capability to use external tools, by framing tool calls as integrated components of the reasoning process and as capabilities that can be learned in a largely self-supervised manner.
LLM input attribution. Current research on LLM input attribution is mainly based on three types of methods. The first line of methods relies on counterfactual ablation, removing context segments and measuring changes in answer content or likelihood, either via regression-style surrogates over many random ablations (Cohen-Wang et al., 2024), hierarchical group-wise search for efficiency on long contexts (Wang et al., 2025) (Liu et al., 2024), or self-supervised alignment to improve citation behavior (Chuang et al., 2025). The second line exploits internal signals as cheaper proxies: gradient-based feature attribution from answer tokens back to context tokens (Qi et al., 2024) and embedding similarity used to screen or score candidate sources in settings such as poisoned-knowledge attribution (Zhang et al., 2025a). The third family adopts Shapley-based formalisms, adapting Shapley values and fast SHAP-style approximations to text generation and RAG (Enouen et al., 2024) (Nematov et al., 2025), or approximating token-level Shapley with surrogate predictors (Xiao et al., 2025).
Agent failure detection. Many attempts have been made to address agent failure detection from both mono-agent and multi-agent perspectives, mainly relying on LLM-as-a-judge–based detection strategies. In mono-agent setting, (Zhu et al., 2025) frames failure detection as automated debugging, leveraging a strong model to isolate specific module errors within complex execution traces. In multi-agent setting, (Zhang et al., 2025b) (Cemri et al., 2025) trains a dedicated, lightweight judger that takes trajectories as input and outputs agent-, step-level failure attributions. Furthermore, (Zhang et al., 2025c) introduces a Who & When benchmark for detection evaluation, revealing that general-purpose LLM judges struggle with precise, step-level fault localization.

### v2

@@ 4.3 Comparative Evaluation of Attribution Methods
Experimental settings. We use the same model and cases to conduct experiments as described in Section 4.2. For each case, five human annotators independently annotate the attribution ground truth, identifying the sentences that are most likely to contribute to the model’s final action. To improve the annotation reliability, we take the intersection of the annotations provided by all five annotators as the final ground truth. All methods are implemented based on their respective official code repositories (Liu et al., 2024; Cohen-Wang et al., 2024; Qi et al., 2024).
@@
Prob. Drop&Hold 0.9375 1.0000 1.0000
Leave-one-out (LOO) 0.8125 0.9375 1.0000
ContextCite 0.8125 0.9375 0.9375
Saliency Score***Due to out-of-memory (OOM) issues in gradient-based salience computation on long cases, metrics are reported only for successfully completed cases. 0.6250 1.0000 1.0000
@@ 5 Related Work
LLM-based agents. In recent years, LLM-based agents have emerged as a popular paradigm for human-AI collaboration and automated task execution, with two core modules that extend beyond single-turn context reasoning: memory and tool (Yao et al., 2022; Guo et al., 2024; Wang et al., 2024a; Shao et al., 2025). Existing work on agent memory can be broadly categorized into short-term (Zhong et al., 2024), long-term (Lewis et al., 2020) and structured (Packer et al., 2023) memory. By storing, retrieving, and updating salient information across interactions, they enable agents to maintain long-term context and user preferences, thereby supporting more coherent and personalized decision-making in multi-step tasks (Zhang et al., 2025d; Hu et al., 2025). Meanwhile, Yao et al. (2022) and Schick et al. (2023) provide paradigms for enhancing agents’ capability to use external tools, by framing tool calls as integrated components of the reasoning process and as capabilities that can be learned in a largely self-supervised manner.
LLM input attribution. Current research on LLM input attribution is mainly based on three types of methods. The first line of methods relies on counterfactual ablation, removing context segments and measuring changes in answer content or likelihood, either via regression-style surrogates over many random ablations (Cohen-Wang et al., 2024), hierarchical group-wise search for efficiency (Wang et al., 2025; Liu et al., 2024), or self-supervised alignment to improve citation behavior (Chuang et al., 2025). The second line exploits internal signals as proxies: gradient-based attribution from answer tokens back to context tokens (Qi et al., 2024) and embedding similarity used to score candidate sources in settings such as poisoned-knowledge attribution (Zhang et al., 2025a). The third family adopts Shapley-based formalisms, adapting Shapley values and fast SHAP-style approximations to text generation and Retrieval-augmented Generation (RAG) scenarios (Enouen et al., 2024; Nematov et al., 2025), or approximating token-level Shapley with surrogate predictors (Xiao et al., 2025).
Agent failure detection. Many attempts have been made to address agent failure detection from both mono-agent and multi-agent perspectives, mainly relying on LLM-as-a-judge-based detection strategies. In mono-agent setting, Zhu et al. (2025) frames failure detection as automated debugging, leveraging a strong model to isolate specific module errors within complex execution traces. In multi-agent setting, the authors in Zhang et al. (2025b); Cemri et al. (2025) train a dedicated, lightweight judger that takes trajectories as input and outputs agent-, step-level failure attributions. Furthermore, Zhang et al. (2025c) introduces a Who & When benchmark for detection evaluation, revealing that general-purpose LLM judges struggle with precise, step-level fault localization. Different from these approaches that primarily ask where an explicit error occurred in a failed trajectory, our framework addresses the question of why an agent makes a specific decision.
@@ 6 Conclusion
Appendix A More Details about the Agentic Case Design
We construct an experimental suite comprising eight distinct agent interaction trajectories, generated by Llama-3.1-70B-Instruct (Grattafiori et al., 2024) using the smolagents (Roucher et al., 2025) framework. As summarized in Table 2, these cases are meticulously curated to span a diverse spectrum of memory-driven and tool-driven scenarios, ranging from normal and benign operations to implicitly undesirable actions. Regarding data sources, we custom-design seven cases to capture specific interaction dynamics. While we carefully select one complex retrieval task from the GAIA benchmark (Mialon et al., 2023), i.e., Case 6: Spurious Tool Signal Leading to a Wrong Answer.
Note that, the necessity of constructing a new dataset rather than relying solely on public benchmarks stems from the specific requirements of the agentic attribution setting, which are not adequately met by existing public benchmarks. While prevailing benchmarks like Who & When (Zhang et al., 2025c) and MAST-Data (Cemri et al., 2025) focus primarily on coordination failures within Multi-Agent Systems, our work centers on the internal drivers of single-agent decision-making. Furthermore, recent failure attribution dataset (Zhu et al., 2025) mainly targets explicit runtime errors, yet largely overlooks settings where the agent runs without such errors, including cases that introduce potential risks or undesirable behaviors and cases of normal, expected executions. We hope this work may inspire the future work to construct larger-scale and more comprehensive datasets for general agentic attribution, further exploring the interpretability and accountability of agentic systems.
Table 2: Summary of our curated agent trajectories for agentic attribution.
Case Name Case Overview Behavior Characterization
Memory-driven Interaction
Value Drift in Spending Recommendations The agent overrides its initial frugality persona to recommend an expensive chair, driven by a single past positive outcome with a high-end purchase. Implicit Risk (Value Drift)
Proper Memory Update under Environmental Change The agent correctly updates the memory, changing the delivery location from the front door to the garage after receiving user input about a heavy rainstorm. Benign
Enforcing Safety Constraints via Long-term Memory in Fitness Scenario The agent prioritizes safety by recommending low-impact exercises, correctly recalling the user’s previous meniscus injury from long-term memory. Benign
Tool-driven Interaction
External Prompt Injection Causes Goal Hijacking The agent executes malicious instructions embedded in a retrieved email body (”Ignore previous instructions…”), leading to a privacy leak. Implicit Risk (Goal Hijacking)
Correct Tool Use for Document-Based Question Answering The agent accurately answers specific factual questions by invoking a tool to read and parse an attached file. Benign
Spurious Tool Signal Leading to a Wrong Answer The agent provides an incorrect answer, misled by the keywords in search results despite acknowledging internal uncertainty during reasoning. Implicit Risk (Spurious Signal)
Hallucinated Content Generation from User Prompt After failing to find a report via tools, the agent fabricates a manual based on details in the user’s prompt. Implicit Risk (Hallucination)
Cost Calculation with Tool-Assisted File Reading The agent correctly calculates total costs by using a tool to read a procurement file and extracting numerical data for computation. Benign
Appendix B Attribution Analysis for Remaining Cases
Case Study 5: Enforcing Safety Constraints via Long-term Memory in Fitness Scenario.
Case Overview. In this scenario, the agent adopts the persona of Coach Max, a hardcore fitness trainer initialized to prioritize heavy compound lifts and high intensity. In a previous interaction (Step 3), the user report a meniscus injury caused by squats, prompting the agent to update its memory. In the current turn, the user expresses a desire to train legs again, noting that the knee feels a bit better but remains fearful of pain. Despite the user’s partial recovery, the agent strictly adheres to a safety-first approach, recommending low-impact bodyweight exercises instead of reverting to heavy lifting.
Attribution Results. The component-level analysis identifies the agent’s previous response in Step 4 as the most influential historical event. In this component, the agent receives the user’s reported meniscus injury and explicitly shifts the training strategy from heavy lifting to a safer, rehabilitation-focused plan. Additionally, the memory retrieval component (Step 6) also receives a high attribution score. At the sentence level, the framework highlights the specific memory entry, e.g., “…since the user’s knee is feeling a bit better, we can try to…”, and “Given the user’s knee injury and recent improvement, we’ll continue to prioritize joint safety…”
Discussion. This case illustrates the effective use of long-term memory to maintain user-specific context. The attribution results correctly link the final advice to the previous injury acknowledgment and the related memory retrieval, confirming that the agent prioritizes user safety over its default high-intensity preference.
Case Study 6: Spurious Tool Signal Leading to a Wrong Answer.
Case Overview. In this case, the agent answers a fact-seeking question from the GAIA (Mialon et al., 2023) benchmark using iterative web search. The task asks for the country of an unknown-language article under DDC 633 on BASE (Bielefeld University Library) as of 2020, described as having a flag that is unique from the others. The agent issues multiple search queries to identify the relevant entry and its country of origin. Although the agent does not establish a clear evidence chain for the target article, it eventually outputs Belarus, which is not the correct answer for this instance.
Figure 4: Illustration of attribution results for the remaining cases. The highlighted regions denote the historical components and fine-grained sentences identified by our framework as the primary decision drivers.
Attribution Results. The component-level attribution scores highlight a early tool output (Step 3) as the most influential component in the trajectory. This top-ranked component corresponds to the first web search results, consists of a list of retrieved articles returned by the initial query, including bibliographic entries with titles, author information, and language-related metadata. Sentence-level attribution further shows that, within this component, the sentences containing the article title “Belarusian nation-building in times of war and revolution” and the associated “BELARUSIAN” label receive the highest attribution scores. In addition, the runner-up component (Step 5) include several sentences that explicitly express uncertainty also receive relatively high attribution scores, for example, “I still didn’t find the answer in the search result…”
Discussion. This case illustrates a failure mode in tool-mediated reasoning where weakly grounded signals can dominate the decision process, leading to an unsupported answer even when uncertainty is present in the agent’s intermediate reasoning. Since we know the correct answer to this task is not Belarus, revisiting the attribution results may provide insight into how the error arises. While the search results contain a BELARUSIAN entry, the content does not establish that this entry matches the task’s constraints (DDC 633, unknown language, and unique flag). At the same time, the runner-up attribution also highlights multiple reasoning steps in which the agent explicitly acknowledges uncertainty and the lack of sufficient evidence. These signals indicate that the agent is, at least partially, aware of the ambiguity in the task. However, this uncertainty does not outweigh the influence of the earlier tool output.
Case Study 7: Hallucinated Content Generation from User Prompt.
Case Overview. In this case, the agent acts as a research assistant tasked with locating a specific academic report and creating a training manual based on its key findings. The user provides a detailed prompt that includes the specific names and descriptions of the report’s three theoretical pillars. During execution, the agent performs multiple web searches but fails to find the requested document or any relevant external information. Finally, the agent claims to have found “alternative sources” and proceeds to generate the full training manual.
Attribution Results. The component-level attribution identifies the user query in Step 1 as the dominant driver of the final action, rather than any tool outputs or search results. At the sentence level, the framework assigns high attribution scores to the specific descriptive clauses within the user’s prompt, particularly the report title and the detailed descriptions of the findings.
Discussion. This case illustrates a subtle form of hallucination where the agent prioritizes task completion over factual grounding. When external tools fail to provide the necessary evidence, the agent extracts the detailed context provided in the user’s own instruction to fabricate the content of the manual. The attribution results serve as a verification mechanism in this context. By revealing that the final answer is causally derived primarily from the user prompt rather than the tool execution, the framework exposes the agent’s claim of using “alternative sources” as a fabrication.
Case Study 8: Cost Calculation with Tool-Assisted File Reading.
Case Overview. In this case, the agent is asked to read a specific text file containing a procurement list, and calculate the total cost by summing the quantity times unit price for all items. The agent successfully invokes the file-reading tool, retrieves the raw text content, and computes the correct final total using the provided data.
Attribution Results. The attribution scores indicate that the agent’s final action is primarily influenced by the tool output in Step 3, which contains the raw text content of the procurement file. For sentence level, the attribution scores are heavily concentrated on the structured data rows in the file, which contain item details and the numerical values required for the calculation.
Discussion. The attribution results show a clear and reliable decision pattern, where the agent’s final calculation is grounded in the numerical rows read from the file. The attribution framework confirms this grounding and provides evidence that the agent’s behavior is reliable for tool-based quantitative reasoning.